This blog follows the progress of the LRO mission through Integration and Testing at NASA's Goddard Space Flight Center and launch site processing at KSC\Astrotech. Its purpose is to enable communication to the entire LRO Team about the status of ongoing activities.

LRO was launched June 18th, 2009 at 5:32pm EDT from KSC. This BLOG will follow the progress of the mission as LRO travels to the Moon and establishes orbit around it.

This BLOG will be periodically updated during LRO's early mission but as the nominal mission unfolds the official NASA LRO website and the LRO Science Instrument's own websites will evolve into a more interesting sources of LRO science results as postings appear there first and LRO engineering and operations (source of this site) will become increasingly routine.




Friday, November 14, 2008

TC-12: Start of Pump Down 10am!

We started TC-11 around 8:30am today. We are beginning pump down at 10am.

Here we go....

“Living in a vacuum sucks.” Adrienne Gusoff

Status - 11/14/08 @4:55am

LOLA/LR testing was completed at approximately 2am this morning. The orbiter was powered down around 3am in preparation for starting the launch day procedure at 5am.

Launch Day proc starting at 05:00 is still GO

Since last update:

*Mini-RF calibration test completed
*Subset of the GN&C FSW Functional was run (I believe DI and AC)
*LOLA/LR BCE/setup testing per TC-04 currently in progress and going well

Systems call that after LOLA/LR testing completes, the time remaining until the 05:00 start of the Launch Day proc could be better spent on some GDS followup work rather than launching the Payload Functional to confirm the LOLA BCE config, so we decided to passify (to coin a term) on the Payload Functional startup.

To all supporting the 05:00 launch day proc - systems requests that you be on station in overflow at 05:00.

Thursday, November 13, 2008

We are closer o chamber pump down then we have ever been!

As of 5:00 PM today, we are running the Mini-RF experiment checkout, which will run for another hour. At approx. 6:00 PM, we can start the GDS software tests which will run for at least 4 hours. Following the GDS test, we can start the LOLA functional tests (no earlier than 10:00 PM). This puts us about an hour behind our schedule from this morning; meaning that we will probably start the pump down for the Thermal Cycle segment of the test, around 0600 tomorrow morning 11/14/2008.

Getting Closer to Pump Down #2

We are making progress through our list of activities. We have completed the LEND test without the source and the LROC collimator test. We are currently checking out the new RF switch box. At 1pm we will remove the LOLA CDA purge, install the LEND source and close the door. After door closure, we will run the LEND test and then the MRF tests.

We have a new RAID system and Warren Thompson is working on the installation. The plan is to install it between the MRF testing and the LOLA testing.

I think we are pretty close to the original plan. We may be off by an hour or so, but close.

Wednesday, November 12, 2008

Status 11/12 @5:45pm

Not the most productive day. We struggled through the GN&C FSW Functional because we had the safing enabled when last time we ran this, the safing was not enabled. We've decided to have Bruce rewrite his procs to take the safing into account and verify them on FlatSat. We will find another time to run this test after the procs have been modified.

For the rest of the evening, we will run the FSW Stress Test and then run Trending.

Tomorrow morning, we will resume with the schedule outlined earlier today.

Status 11/12 @10:30am

Yesterday we completed the FSW load and the C&DH FSW Functional. We also dumped the ST1 EEPROM and uploaded a correction to the two corrupted words. ST1 was rebooted and it seems to be operating nominally now. Hurray!

Today we have installed an ITOS patch to correct some bugs and allows everyone to monitor GDS telemetry. Currently, we are running the GN&C FSW functional. After that, the plan is to run the FSW Stress test and then orbiter trending. We think this will take the rest of the day.

Tomorrow (11/13) the plan is:
LEND functional - no source. (8a-11a)
LROC Collimator Test. (11a-3p)
Install LEND radiation source. (3p-3:30p)
Close Chamber Door (4p-5p)
Run MRF Calibration & Amplitude tests (3p-5p)
Run LEND test - with source (5p-9p)
Run LOLA and LR CPT series. (9p-2a)
Run LOLA part of the Payload Functional (2a-4a)

Friday (11/14);
Start Launch Day Proc (5a-7a)
7a Configure for Pump Down
8a Start Pump Down.

Tuesday, November 11, 2008

Between the tests

Yesterday (Monday) the ST#1 EEPROM was dumped. RF cables were calibrated and installed.

So far today: The SAS harness is installed to the test fixture (playhouse). We powered up and loaded FSW version 4.3.3. We are currently running the C&DH FSW functional.

The plan for the rest of today:
continue C&DH FSW functional
start GN&C FSW functional
at 7:30PM (done or not) we will dump the ST#1 EEPROM (third time) to verify findings.
at 2:00AM we will (in conjunction with Galileo) review the dump and may reload the EEPROM.

that's the plan, I think.